# Models for rational $(\infty,1)$-categories

### v1

@@ 1. Introduction
The study of $(\infty,1)$-categories is intended to understand higher categorical structures that are equipped with a notion of homotopy between their morphisms. Namely, an $(\infty,1)$-category is a categorical structure that is enriched in spaces. Thus, between any two objects, we have a space of $1$-morphisms; a homotopy, or $2$-morphism, is a path in that mapping space. Moreover, the enrichment in spaces yields homotopies between homotopies and so on, so we have morphisms in infinitely high degrees. However, what gives us control over this structure is that all morphisms in degrees higher than $1$ are invertible up to homotopy because homotopies can be traveled in reverse.
@@
Our understanding of $(\infty,1)$-categories has been advanced thanks to the development of various models for $(\infty,1)$-categories, that is, mathematical objects that exhibit the structure of an $(\infty,1)$-category. More specifically, Quillen’s concept of a model category from [Qui67] constitutes a natural and convenient framework for the purposes of modeling $(\infty,1)$-categories. Conceptually, a model structure on a category packages a homotopy theory on that category, for it allows us to define homotopies, function complexes, and an associated homotopy category. In this framework, one not only introduces mathematical objects that encode the structure of an $(\infty,1)$-category, but also develops model categories that describe the homotopy theory of $(\infty,1)$-categories.
@@
There is a wide variety of equivalent models for $(\infty,1)$-categories; a treatment of those models and their equivalence is provided by Bergner in [Ber18]. In this paper, we use two such models, both of which are particular kinds of simplicial spaces: complete Segal spaces and Segal categories. Complete Segal spaces, as defined by Rezk in [Rez01], are simplicial spaces satisfying a condition that encodes morphism composition up to homotopy. Rezk’s work yields a cartesian model category whose fibrant objects are precisely the complete Segal spaces; this model category is produced as a left Bousfield localization of the Reedy model structure on simplicial spaces.
@@
Segal categories are also simplicial spaces that satisfy the same condition and have a discrete space of 0-simplices; the discreteness condition accounts for the fact that we encode a set of objects. Segal categories were introduced by Dwyer, Kan, and Smith in [DKS89], albeit under a different name, and they were developed by Hirschowitz and Simpson in [HS], as well as by Pellissier in [Pel] from the homotopical perspective. In [Ber07c], Bergner develops two equivalent model categories for the purposes of studying Segal categories, and she establishes that both model categories are equivalent to Rezk’s model category for complete Segal spaces.
@@
Having such a rich theory at their disposal, one may be motivated to investigate $(\infty,1)$-categories enriched in a class of spaces with a prescribed homotopical property; a goal of such an investigation can be to define equivalent analogs of the existing models for $(\infty,1)$-categories. For instance, in the setting of equivariant homotopy theory, equivariant complete Segal spaces were defined by Bergner and Chadwick in [BC15], and Bergner also produced equivalent models for $(\infty,1)$-categories enriched in spaces with a group action in [Ber17].
@@
In this paper, we provide analogs of complete Segal spaces and Segal categories for rational homotopy theory, which we then show to be equivalent. In order not to require mapping spaces to be simply connected, we adopt the framework for the rational homotopy theory of non-simply connected spaces due to Gómez-Tato, Halperin, and Tanré [GTHT00]. In that framework, a space is said to be rational if its higher homotopy groups are rational vector spaces, with no conditions imposed on its fundamental group or its set of path components.
@@
With this notion of a rational space, we provide two equivalent models for rational $(\infty,1)$-categories, which we define to be $(\infty,1)$-categories enriched in rational spaces. Our first model is that of rational complete Segal spaces, which are complete Segal spaces with rational mapping spaces, and our second model is that of rational Segal categories, which are Segal categories with rational mapping spaces. By a left Bousfield localization of Rezk’s model category for complete Segal spaces and Bergner’s model category for Segal categories, we produce a model category for rational complete Segal spaces and a model category for rational Segal categories, respectively. We also establish that our model category for rational complete Segal spaces is cartesian; to do so, we apply Rezk’s criterion from [Rez01] on when a localization of the Reedy model structure remains compatible with the cartesian closure of simplicial spaces. Moreover, the equivalence between complete Segal spaces and Segal categories descends to an equivalence between rational complete Segal spaces and rational Segal categories.
@@
There is a model category whose fibrant objects are the rational complete Segal spaces. There is also a model category whose fibrant objects are the Reedy fibrant rational Segal categories. These two model categories are Quillen equivalent.
@@
We also investigate the relationship between complete Segal spaces and Segal categories through the lens of the left transfer of a model category, as developed by Bayeh, Hess, Karpova, Kędziorek, Riehl, and Shipley in [BHK+15]. In particular, we show that the model category for Segal categories can be described as a left transfer of the model category for complete Segal spaces, as well as that our model category for rational Segal categories can be described as a left transfer of our model category for rational complete Segal spaces. The former left transfer might be of independent interest as it greatly simplifies Bergner’s proof from [Ber07c].
The techniques that we employ in this paper are also intended to yield rational analogs of other known models for $(\infty,1)$-categories. Such models are quasi-categories, as introduced by Boardman and Vogt in [BV73] and studied by Joyal in [Joy02] and Lurie in [Lur09], as well as simplicial categories, as studied by Bergner in [Ber07b]. We are working on introducing rational analogs of these two models, and then showing that they are equivalent to the two equivalent models for rational $(\infty,1)$-categories that we introducing in this paper.
Lastly, we remark that we develop our theory in greater generality than that maintained thus far in this introduction. For a multiplicative subset $M$ of $\mathbb{Z}$, we say that a space is $M$-local if its higher homotopy groups are $M^{-1}\mathbb{Z}$-vector spaces. Choosing $M=\mathbb{Z}\setminus\{0\}$ recovers the rational case that we have focused on, but this higher level of abstraction grants one the flexibility of localizing at a specific set of prime numbers. We give two equivalent models for $(\infty,1)$-categories enriched in $M$-local spaces: our first model is that of $M$-local complete Segal spaces, which are complete Segal spaces that with $M$-local mapping spaces, and our second model is that of $M$-local Segal categories, which are Segal categories with $M$-local mapping spaces.
@@
Section 2 comprises the necessary background in simplicial sets and model categories. In Section 3, we establish our characterization of rational spaces in terms of function complexes. In Section 4, we use this characterization to give our model category for rational spaces. Section 5 has background on complete Segal spaces. Section 6 is on Segal categories, and it contains the description of the model category for Segal categories as a left transfer of the model category for complete Segal spaces. In Section 7, we introduce rational Segal categories and we produce a model category for them. In Section 8, we introduce rational complete Segal spaces and we produce a cartesian model category for them. Lastly, in Section 9, we establish the equivalence of rational complete Segal spaces and rational Segal categories, and we describe our model category for rational Segal categories as a left transfer of our model category for rational complete Segal spaces.
@@ 2.1. Simplicial sets
Simplicial sets are a combinatorial model for spaces; our reference is [GJ99, Chapter I].
@@
We write $\mathcal{SS}\mathrm{ets}$} for the category of simplicial sets and simplicial maps (natural transformations) between them.
@@ 2.2. Model categories
We give a brief review of model categories from [DS95, §3] and [Hir03, Chapters 7-13].
@@
A model category comprises a category $\mathcal{M}$ that has all finite limits and colimits, as well as three distinguished classes of morphisms in $\mathcal{M}$: a class of weak equivalences, a class of cofibrations, and a class of fibrations. A weak equivalence that is also a cofibration is called an acyclic cofibration. A weak equivalence that is also a fibration is called an acyclic fibration. This data satisfies the axioms in [DS95, Definition 3.3].
@@
    a weak equivalence if and only if $f$ is a weak homotopy equivalence;
@@
The model category $\mathcal{SS}\mathrm{ets}$} has a list of features that we rely on, but do not explicitly use in this paper; namely, $\mathcal{SS}\mathrm{ets}$} is combinatorial [Dug01], left proper [Hir03, Corollary 13.1.4], and simplicial [Hir03, Example 9.1.13]. In particular, for two simplicial sets $X$ and $Y$, we have a function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(X,Y)$ with $n$-simplices
@@
Theorem 2.11 ([GJ99, Chapter I, Theorem 11.4]).
@@ 2.3. Left Bousfield localization
We give an overview of left Bousfield localization from [Hir03, Chapter 4]. Let $\mathcal{M}$ be a simplicial model category; in particular, for any two objects $X$ and $Y$ of $\mathcal{M}$, we have a function complex $\operatorname{Map}_{\mathcal{M}}(X,Y)$. Let $T$ be a set of morphisms in $\mathcal{M}$.
@@
    An object $K$ of $\mathcal{M}$ is $T$-local if $K$ is fibrant in $\mathcal{M}$ and, for every morphism $\psi\colon A\rightarrow B$ in $T$, the map $\operatorname{Map}_{\mathcal{M}}(\psi,K)\colon\operatorname{Map}_{\mathcal{M}}(B,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(A,K)$ is a weak homotopy equivalence.
@@
    A morphism $f\colon X\rightarrow Y$ in $\mathcal{M}$ is a $T$-local equivalence if, for every $T$-local object $K$ of $\mathcal{M}$, the map $\operatorname{Map}_{\mathcal{M}}(f,K)\colon\operatorname{Map}_{\mathcal{M}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(X,K)$ is a weak homotopy equivalence.
@@ 3. Rational homotopy theory of non-simply connected spaces
From now on, let $M$ be a multiplicative subset of $\mathbb{Z}$. We gather some background on the rational homotopy theory of non-simply connected spaces. In particular, we focus on the spaces whose higher homotopy groups are $M^{-1}\mathbb{Z}$-modules; we refer to such spaces as $M$-local spaces. When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we arrive at the study of rational spaces. Our main result in this section is our characterization of $M$-local spaces in terms of function complexes, which is a cornerstone for our arguments in this paper. Although we do not expect this result to be novel to the rational homotopy theory expert, we have not found it in the literature, so we provide a self-contained proof.
@@ 3.1. The construction of local spheres
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. Our work in this section makes reference to the CW complex construction of the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ from [FHT01, Chapter 9(a)], which subsumes the construction of the rational $n$-sphere $S^{n}_{\mathbb{Q}}$. We go over this construction.
@@ 3.2. A characterization of rational spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak homotopy equivalence.
@@
If $Y$ is $M$-local, then Proposition 3.3 implies that $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak homotopy equivalence for every $n\geq 2$. Conversely, for $n\geq 2$, let $[f]$ be an element of $\pi_{n}(Y)$ represented by a map $f\colon S^{n}\rightarrow Y$, and let $m\in M$. The induced bijection of path components $S^{n}_{M^{-1}\mathbb{Z}},Y]\xrightarrow{\cong}[S^{n},Y]$ informs us that there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
@@
4. The model category for rational spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Using our characterization from Section 3, we left Bousfield localize the model category of simplicial sets to get a model category whose fibrant objects are the $M$-local Kan complexes. In particular, we obtain a model category whose fibrant objects are the rational Kan complexes.
Our characterization of $M$-local spaces in Proposition 3.4 prompts us to left Bousfield localize $\mathcal{SS}\mathrm{ets}$} at the set

### v2

@@ Abstract.
Key words and phrases:
$(\infty,1)$-categories, rational homotopy theory, model categories, complete Segal spaces, Segal categories.
1991 Mathematics Subject Classification:
55U35 (Primary), 18N60, 55P62, 18N55, 55P60, 18N50, 55U10, 18N40 (Secondary).
@@ 1. Introduction
The study of $(\infty,1)$-categories is intended to understand categories with a notion of homotopy between their morphisms. Namely, an $(\infty,1)$-category is a category enriched in spaces, possibly weakly. This enrichment yields $k$-morphisms for all $k$ given by homotopies. However, all morphisms in dimension higher than $1$ are invertible up to homotopy because homotopies can be traveled in reverse.
@@
Our understanding of $(\infty,1)$-categories has been advanced thanks to the development of various models for $(\infty,1)$-categories, that is, mathematical objects that exhibit the structure of an $(\infty,1)$-category. More specifically, Quillen’s concept of a model category from [23] constitutes a natural and convenient framework for the purpose of modeling $(\infty,1)$-categories. Conceptually, a model structure on a category packages a homotopy theory on that category, for it allows us to define homotopies, function complexes, and an associated homotopy category. In this framework, one not only introduces mathematical objects that encode the structure of an $(\infty,1)$-category, but also develops model categories that describe the homotopy theory of $(\infty,1)$-categories.
@@
There is a wide variety of equivalent models for $(\infty,1)$-categories; a treatment of those models and their equivalence is provided by Bergner in [7]. In this paper, we use two such models, both of which are particular kinds of simplicial spaces: complete Segal spaces and Segal categories. Complete Segal spaces, as defined by Rezk in [25], are simplicial spaces satisfying a condition that encodes morphism composition up to homotopy. Rezk’s work yields a cartesian model category whose fibrant objects are precisely the complete Segal spaces; this model category is produced as a left Bousfield localization of the Reedy model structure on simplicial spaces.
@@
Segal categories are also simplicial spaces that satisfy the same condition and have a discrete space of 0-simplices; the discreteness condition accounts for the fact that we encode a set of objects. Segal categories were introduced by Dwyer, Kan, and Smith in [12], albeit under a different name, and they were developed by Hirschowitz and Simpson in [19], as well as by Pellissier in [22] from the homotopical perspective. In [5], Bergner develops two equivalent model categories for the purpose of studying Segal categories, and she establishes that both model categories are equivalent to Rezk’s model category for complete Segal spaces.
@@
Having such a rich theory at their disposal, one may be motivated to investigate $(\infty,1)$-categories enriched in a class of spaces with a prescribed homotopical property; a goal of such an investigation can be to define equivalent analogs of the existing models for $(\infty,1)$-categories. For instance, in the setting of equivariant homotopy theory, equivariant complete Segal spaces were defined by Bergner and Chadwick in [8], and Bergner also produced equivalent models for $(\infty,1)$-categories enriched in spaces with a group action in [6].
@@
In this paper, we provide analogs of complete Segal spaces and Segal categories for rational homotopy theory, which we then show to be equivalent. In order not to require mapping spaces to be simply connected, we adopt the framework for the rational homotopy theory of non-simply connected spaces due to Gómez-Tato, Halperin, and Tanré [16]. In that framework, a space is said to be rational if its higher homotopy groups are rational vector spaces, with no conditions imposed on its fundamental group or its set of path components.
@@
With this notion of a rational space, we provide two equivalent models for rational $(\infty,1)$-categories, which we define to be $(\infty,1)$-categories enriched in rational spaces. Our first model is that of rational complete Segal spaces, which are complete Segal spaces with rational mapping spaces, and our second model is that of rational Segal categories, which are Segal categories with rational mapping spaces. By a left Bousfield localization of Rezk’s model category for complete Segal spaces and Bergner’s model categories for Segal categories, we produce a model category for rational complete Segal spaces and two model categories for rational Segal categories, respectively. We also establish that our model category for rational complete Segal spaces is cartesian; to do so, we apply Rezk’s criterion from [25] on when a localization of the Reedy model structure remains compatible with the cartesian closure of simplicial spaces. Moreover, the equivalence between complete Segal spaces and Segal categories descends to an equivalence between rational complete Segal spaces and rational Segal categories.
@@
There exist:
  1. (1)
    a model category whose fibrant objects are the rational complete Segal spaces;
  2. (2)
    a model category whose fibrant objects are the Reedy fibrant rational Segal categories; and
  3. (3)
    a model category whose fibrant objects are the levelwise fibrant rational Segal categories.
These model categories are Quillen equivalent.
@@
In fact, our argument is not exclusive to rational spaces, and it works for any class of spaces that arises from a left Bousfield localization of the model structure on spaces. For example, in the chromatic homotopy-theoretic framework developed by Heuts in [17], our work yields two equivalent models for $(\infty,1)$-categories enriched in $v_{n}$-periodic spaces for a non-negative integer $n$, with the case $n=0$ being that of rational $(\infty,1)$-categories.
In the process, we investigate the relationship between complete Segal spaces and Segal categories through the lens of the left transfer of a model category, as developed by Bayeh, Hess, Karpova, Kędziorek, Riehl, and Shipley in [2]. We show that the model category for Reedy fibrant Segal categories can be described as a left transfer of the model category for complete Segal spaces, as well as that our model category for Reedy fibrant rational Segal categories is left-induced from our model category for rational complete Segal spaces. The former left transfer might be of independent interest as it greatly simplifies Bergner’s proof from [5].
Lastly, we hope that the techniques in this paper help us develop rational homotopy-theoretic analogs of other models for $(\infty,1)$-categories in the future. Such models are quasi-categories, as introduced by Boardman and Vogt in [9] and studied by Joyal in [20] and Lurie in [21], as well as simplicial categories, as studied by Bergner in [4]. We are working on introducing rational analogs of these two models, and then showing that they are equivalent to the two equivalent models for rational $(\infty,1)$-categories that we introduce in this paper.
@@
Section 2 comprises the necessary background in simplicial sets and model categories. In Section 3, we establish our characterization of rational spaces in terms of function complexes, and we use it to give our model category for rational spaces. Section 4 has background on complete Segal spaces. Section 5 is on Segal categories, and it contains the description of the model category for Reedy fibrant Segal categories as a left transfer of the model category for complete Segal spaces. In Section 6, we introduce rational Segal categories and we give two model categories for them. In Section 7, we introduce rational complete Segal spaces and we produce a cartesian model category for them. Lastly, in Section 8, we establish the equivalence of rational complete Segal spaces and rational Segal categories, and we describe our model category for Reedy fibrant rational Segal categories as a left transfer of our model category for rational complete Segal spaces.
@@ 2.1. Simplicial sets
Simplicial sets are a combinatorial model for spaces; our reference is [15, Chapter I].
@@
We write $\mathcal{SS}\mathrm{ets}$} for the category of simplicial sets and their simplicial maps (natural transformations).
@@ 2.2. Model categories
We give a brief review of model categories from [13, §3] and [18, Chapters 7-13].
@@
A model category comprises a category $\mathcal{M}$ that has all finite limits and colimits, as well as three distinguished classes of morphisms in $\mathcal{M}$: a class of weak equivalences, a class of cofibrations, and a class of fibrations. A weak equivalence that is also a cofibration is called an acyclic cofibration. A weak equivalence that is also a fibration is called an acyclic fibration. This data satisfies the axioms in [13, Definition 3.3].
@@
    a weak equivalence if and only if $|f|$ is a weak homotopy equivalence;
@@
The model category $\mathcal{SS}\mathrm{ets}$} has a list of features that we rely on, but do not explicitly use in this paper; namely, $\mathcal{SS}\mathrm{ets}$} is combinatorial [10], left proper [18, Corollary 13.1.4], and simplicial [18, Example 9.1.13]. In particular, for two simplicial sets $X$ and $Y$, we have a function complex $\operatorname{Map}_{{\mathcal{SS}\mathrm{ets}}}(X,Y)$ with $n$-simplices
@@
Theorem 2.11 ([15, Chapter I, Theorem 11.4]).
@@ 2.3. Left Bousfield localization
We give an overview of left Bousfield localization from [18, Chapter 4]. Let $\mathcal{M}$ be a simplicial model category; in particular, for any two objects $X$ and $Y$ of $\mathcal{M}$, we have a function complex $\operatorname{Map}_{\mathcal{M}}(X,Y)$. Let $T$ be a set of morphisms in $\mathcal{M}$.
@@
    An object $K$ of $\mathcal{M}$ is $T$-local if $K$ is fibrant in $\mathcal{M}$ and, for every morphism $\lambda\colon A\rightarrow B$ in $T$, the map $\operatorname{Map}_{\mathcal{M}}(\lambda,K)\colon\operatorname{Map}_{\mathcal{M}}(B,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(A,K)$ is a weak equivalence.
@@
    A morphism $f\colon X\rightarrow Y$ in $\mathcal{M}$ is a $T$-local equivalence if, for every $T$-local object $K$ of $\mathcal{M}$, the map $\operatorname{Map}_{\mathcal{M}}(f,K)\colon\operatorname{Map}_{\mathcal{M}}(Y,K)\rightarrow\operatorname{Map}_{\mathcal{M}}(X,K)$ is a weak equivalence.
@@
2.4. Cartesian model categories
Lastly, we go over cartesian model categories following [7, Chapter 2.9].
Definition 2.15.
A category $\mathcal{C}$ is cartesian closed if it has finite products and, for any two objects $X$ and $Y$, an internal function object $Y^{X}$ satisfying, for any object $Z$, a natural isomorphism
$\operatorname{Hom}_{\mathcal{C}}(Z,Y^{X})\cong\operatorname{Hom}_{\mathcal{C}}(Z\times X,Y).$
A cartesian model category is a cartesian closed category with a compatible model structure.
Definition 2.16 ([26, §2.2]).
A model category $\mathcal{M}$ is cartesian if its underlying category is cartesian closed, its terminal object is cofibrant, and, given a cofibration $i\colon A\rightarrow A^{\prime}$ and a fibration $p\colon X^{\prime}\rightarrow X$, the induced map
$q\colon(X^{\prime})^{A^{\prime}}\rightarrow(X^{\prime})^{A}\underset{X^{A}}{\times}X^{A^{\prime}}$
is a fibration that is acyclic if $i$ or $p$ is acyclic.
The model category $\mathcal{SS}\mathrm{ets}$} is cartesian, for its cartesian structure coincides with its simplicial structure.
@@ 3. Rational homotopy theory of non-simply connected spaces
From now on, let $M$ be a multiplicative subset of $\mathbb{Z}$. We study the rational homotopy theory of non-simply connected spaces. In particular, we focus on the spaces whose higher homotopy groups are $M^{-1}\mathbb{Z}$-modules; we refer to such spaces as $M$-local spaces. When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we arrive at the study of rational spaces. Our main result is a characterization of $M$-local spaces in terms of function complexes, which we use to left Bousfield localize the model category of simplicial sets and get a model category whose fibrant objects are the $M$-local Kan complexes. In particular, we obtain a model category whose fibrant objects are the rational Kan complexes.
@@ 3.1. The construction of local spheres
Let $M$ be a multiplicative subset of $\mathbb{Z}$ and let $n\geq 2$. Our work in this section makes reference to the CW complex construction of the $M$-local $n$-sphere $S^{n}_{M^{-1}\mathbb{Z}}$ from [14, Chapter 9(a)], which subsumes the construction of the rational $n$-sphere $S^{n}_{\mathbb{Q}}$. We go over this construction.
@@ 3.2. A characterization of rational spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A space $Y$ is $M$-local if and only if, for every $n\geq 2$, the restriction map $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence.
@@
If $Y$ is $M$-local, then Proposition 3.3 implies that $\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n}_{M^{-1}\mathbb{Z}},Y)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}}(S^{n},Y)$ is a weak equivalence for every $n\geq 2$. Conversely, for $n\geq 2$, let $[f]$ be an element of $\pi_{n}(Y)$ represented by a map $f\colon S^{n}\rightarrow Y$, and let $m\in M$. The induced bijection of path components $S^{n}_{M^{-1}\mathbb{Z}},Y]\xrightarrow{\cong}[S^{n},Y]$ informs us that there exists a map $g\colon S^{n}_{M^{-1}\mathbb{Z}}\rightarrow Y$, unique up to homotopy, such that the diagram
@@
3.3. The model category for rational spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Our characterization of $M$-local spaces in Proposition 3.4 prompts us to left Bousfield localize $\mathcal{SS}\mathrm{ets}$} at the set